OpenStudy (teller):

Can someone help me please to see if I did this correctly ? Suppose that the functions r and s are defined for all real numbers x as follows. r(x)=2x+2 s(x)=x-2 Write the expressions for (s-r)(x) and (s+r)(x) and evaluate (s*r)(-2) (s-r)(x)= -x-7 (s+r)(x)=3x-3) (s*r)(-2)= I'm having a little trouble on the last part on how to solve it.

jimthompson5910 (jim_thompson5910):

how did you get `(s-r)(x)= -x-7` ?

OpenStudy (teller):

I kept getting 61 but it said it was wrong I placed it in my calculator and I know that's right because it told me I was write for -x-7 but I seem to be getting the last part wrong.

OpenStudy (teller):

right not write

jimthompson5910 (jim_thompson5910):

x-2x = -x, that is correct -2-2 = -7 is not correct

OpenStudy (teller):

But it told me I was correct....

OpenStudy (teller):

the only thing it told me I was wrong on was the last part ...

jimthompson5910 (jim_thompson5910):

there has to be a typo somewhere

OpenStudy (teller):

I don't see it I copied and paste it want me to send you the screen shot?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (teller):

This is what it says

jimthompson5910 (jim_thompson5910):

I see, it should be s(x) = x-5

OpenStudy (teller):

Nope because before I put that and it told me that they need to be simplified further and that's what I got when I simplified it ...

jimthompson5910 (jim_thompson5910):

No I meant the original problem

jimthompson5910 (jim_thompson5910):

instead of s(x) = x-2 the actual problem is s(x) = x-5

OpenStudy (teller):

So it should be this ? x-5*(2x+2)?

OpenStudy (teller):

Then I would solve it that way?

jimthompson5910 (jim_thompson5910):

(s*r)(-2)= s(-2)*r(-2) what are the values of s(-2) and r(-2) ?

OpenStudy (teller):

Um I'm not sure...

jimthompson5910 (jim_thompson5910):

s(x) = x-5 s(-2) = -2 - 5 ... replace every x with -2 s(-2) = ??

OpenStudy (teller):

-2-5*(2(-2)+2)?

jimthompson5910 (jim_thompson5910):

you need parenthesis around the `-2-5`

OpenStudy (teller):

so (-2-5)*(2(-2)+2)?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (teller):

14?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (teller):

Alright thank you!

jimthompson5910 (jim_thompson5910):

s(x) = x-5 s(-2) = -2 - 5 ... replace every x with -2 s(-2) = -7 r(x) = 2x+2 r(-2) = 2(-2)+2 r(-2) = -2 (s*r)(-2)= s(-2)*r(-2) (s*r)(-2)= -7*(-2) (s*r)(-2)= 14

OpenStudy (teller):

Thanks!